Determine the area under the standard normal curve that lies to the right of

left parenthesis a right parenthesis Upper Z equals 1.25 comma(a) Z=1.25,

​(b)

Upper Z equals negative 0.86 commaZ=−0.86,

​(c)

Upper Z equals 0.02 commaZ=0.02,

and​ (d) Upper Z equals negative 0.78 .Z=−0.78.
Answer #1

Answer

We have to find the area to the rigth of given z values

(A) Using excel function 1- NORMSDIST(z) for right area

setting z = 1.25

this implies

= 1- NORMSDIST(1.25)

= 1- 0.8944

= 0.1056

(B) Using excel function 1- NORMSDIST(z) for right area

setting z = -0.86

this implies

= 1- NORMSDIST(-0.86)

= 1- 0.1949

= 0.8051

(C) Using excel function 1- NORMSDIST(z) for right area

setting z = 0.02

this implies

= 1- NORMSDIST(0.02)

= 1- 0.5080

= 0.4920

(D) Using excel function 1- NORMSDIST(z) for right area

setting z = -0.78

this implies

= 1- NORMSDIST(-0.78)

= 1- 0.2177

= 0.7823